### v1

# URL: ]https://photox.vscht.cz/URL: ]https://www.controlled-molecule-imaging.org

Abstract
Aqueous radiation chemistry emerges through ultrafast proton transfer and ion-radical formation with unexplored energy-redistribution dynamics steering the subsequent reactions. We performed time-resolved disruptive probing on pure water dimer, $(\text{H}_{2}\text{O})_{2}$, to disentangle the post-ionization reactions. Through kinetic-energy-resolved ion imaging, we unraveled the dynamics in the $(\text{H}_{2}\text{O})_{2}^{+}$ ground state: at low-energy ($\mathord{\sim}$0.05 eV) ultrafast proton transfer ($\mathord{\sim}$19 fs) is followed by $\text{H}_{3}\text{O}^{+}+\text{OH}$ fragmentation ($\mathord{\sim}$360 fs). At higher energies, proton transfer becomes hindered ($\mathord{\sim}$60 fs) while the subsequent fragmentation becomes faster ($\mathord{\sim}$210 fs), evolving into coupled dynamics (>0.15>0.15 eV, $\mathord{\sim}$100 fs). Moreover, we observed $(\text{H}_{2}\text{O})_{2}^{+}$ stabilization proceeding through a Zundel-like structure. This reveals how ion-radical formation in ionized hydrogen-bonded networks shapes reactivity in aqueous dynamics.
Radiation chemistry in aqueous environments is largely governed by water radiolysis [1, 2, 3]. Upon irradiation, water undergoes a cascade of reactions initiated by the formation of highly reactive ion–radical species [3, 4] that give rise to fundamental phenomena such as biological radiation damage, catalytic transformations, material corrosion, and radical-driven oxidation. A molecular-level characterization of transient reactive species formation and the associated energy redistribution is therefore essential for advancing technologies including radiation therapy [4, 5], wastewater treatment [6, 7], nuclear reactor coolant systems [4, 8], and spaceflight applications [9, 10].
Surprisingly, the initial elementary steps driving the subsequent post-ionization dynamics in water were only recently explored in a time-resolved fashion [3, 2, 11]. These studies provide a coarse picture lacking state- and energy-resolved details. The dominant initial step following valence ionization of liquid water is the ultrafast proton transfer (PT) reaction occurring on a sub-50-fs timescales [12, 3, 13], producing a hydronium cation ($\text{H}_{3}\text{O}^{+}$) and a hydroxyl radical (OH):
$\text{H}_{2}\text{O}\cdots{}\text{H}_{2}\text{O}\xrightarrow[\text{ionization}]{-e^{-}_{\text{valence}}}\text{H}_{2}\text{O}^{+}{}\cdots\text{H}_{2}\text{O}\xrightarrow{\text{PT}}\text{H}_{3}\text{O}^{+}{}\cdots{}\text{OH},$ (1)
These products drive secondary PTs, oxidative reactions, or undergo recombination. Early pump-probe studies lacked the temporal resolution to capture the initial PT dynamics [12]. Time-resolved x-ray absorption spectroscopy (XAS), enabled by x-ray free-electron lasers, revealed an ultrafast $\sim$46 fs PT event [3], while ultrafast electron diffraction captured $\text{H}_{3}\text{O}^{+}\cdotsOH$ formation in $\mathord{\sim}140$ fs followed by the ion–radical separation occurring on a $\sim$250 fs timescale [14]. These seemingly different timescales reflect the complementary sensitivities of the probes: XAS probes local electronic structure and diffraction probes atomic motion. Recently, experiments on water-dimer clusters ionized by 24 eV photons reported a PT time of $\sim$55 fs, consistent with the earlier XAS results [2].
Molecular clusters permit direct detection of ionic fragments, providing detailed insights into energy redistribution along ultrafast reaction pathways [15], whereas secondary processes characteristic of the bulk liquid [16] remain inaccessible. We used water dimer, $(\text{H}_{2}\text{O})_{2}$, a controllable and well-defined model [17, 18, 19], to investigate post-ionization chemistry in aqueous systems.
Figure 1: Schematic picture of the experiment. Water dimers were strong-field ionized to launch the ultrafast dynamics. The dynamics were perturbed by a delayed, weak probe pulse that redistributed population and modified the fragmentation pattern. Monitoring all ion species over the pump–probe delay provided time-resolved signatures of PT, ion–radical stabilization, or fragmentation. Velocity-map imaging further yielded delay-dependent momentum distributions that reveal the interplay between PT and fragmentation.
Previously, we found a surprising diversity of ion products after single ionization of $(\text{H}_{2}\text{O})_{2}$. Our approach using the electrostatic deflector [20, 17] allows the assignment of the originally ionized system without the need for simultaneous detection of multiple ions, as in coincidence-detection methods [2]. Similar to earlier studies [3, 11], we populated $(\text{H}_{2}\text{O})_{2}^{+}$ via strong-field ionization, which raises the question of the initial electronic states. Our recent study showed that ionization predominantly accesses the four energetically lowest states, $\text{D}_{0}$$\text{D}_{3}$ [21]. According to our previous ab initio simulations [2] these undergo PT, forming the ion–radical complex ($\text{H}_{3}\text{O}\cdots\text{OH})^{+}$, which either stabilizes, predominantly for ionization into $\text{D}_{0}$, or fragments to $\text{H}_{3}\text{O}^{+}+\text{OH}$.
To explore the elementary dynamics of ionized water dimer, we employed the mass-spectrometric disruptive-probing (DP) method. DP was recently presented as a panoramic approach [22] that synchronously monitors delay-dependent signal changes arising from perturbations of evolving electronic and nuclear wavepackets in multiple independent photochemical pathways [23]. A schematic overview of the approach is shown in Figure 1 : A strong-field ($I_{\text{pump}}=2.3\cdot 10^{14}~\text{W}/\text{cm}^{2}$, 800800 nm) pump pulse initiates the dynamics, while a weaker, delayed NIR-probe pulse perturbs the ongoing evolution. When temporally separated from the pump, the probe is too weak to induce ionization. Regardless of the precise disruption mechanism [22, 23], the probe modifies the yields of individual ion species as the delay is scanned, thereby encoding the reaction dynamics. So far, DP experiments only took ion yields into account [22]. Our combination of DP with velocity-map imaging (VMI) adds kinetic-energy resolution to the transient ion signals. Complementing these experiments with ab initio simulations of the dynamics to 1 ps allowed us to assign fragmentation pathways beyond the initial PT event.
We disentangled the ultrafast dynamics of $(\text{H}_{2}\text{O})_{2}^{+}$ in its electronic ground state, determined energy-resolved timescales for PT ($20\ldots 100$ fs), fragmentation ($360\ldots 100$ fs), and their coalescence at higher energies as well as $(\text{H}_{2}\text{O})_{2}^{+}$ stabilization ($\mathord{\sim}1$ ps). This directly provides the energy redistribution in the hydronium $\text{H}_{3}\text{O}^{+}$, OH radical, and ion-radical formation in ionized aqueous systems.
Results and discussion
Figure 2: Recorded transient signals related to disruptive probing of the $\text{H}_{3}\text{O}^{+}$ channel and their analysis. A and B display transient signals of the $\text{H}_{3}\text{O}^{+}$ and $\text{H}_{2}\text{O}^{+}$ channels (blue dots and black error bars), respectively. Fits (red lines) were obtained using the parametric model described in Materials and Methods. The data (gray points) in the range (500,50)(-500,-50) fs were excluded from the fitting procedure due to probe-induced pre-excitation effects. The inset in A shows a velocity map of $\text{H}_{3}\text{O}^{+}$ with only the central part (yellow circle) originating from $(\text{H}_{2}\text{O})_{2}^{+}$ dynamics. C displays the result of simulations of $(\text{H}_{2}\text{O})_{2}^{+}$ dynamics in its ground electronic state ($\text{D}_{0}$), corresponding data for $\text{D}_{1}$$\text{D}_{3}$ are provided in Supplementary Text. D represents the background corrected delay- and energy-resolved $\text{H}_{3}\text{O}^{+}$ signal. E shows the comparison of the bleached (blue), pump-only (yellow), and simulated ($\text{D}_{0}$, red) kinetic energy distributions [21]. F displays the timescales of the $\text{H}_{3}\text{O}^{+}$ formation obtained from the kinetic energy-resolved analysis using the parametric models described in Materials and Methods; see also Supplementary Text for the individual fits.
Figure 3: Recorded transient signals related to disruptive probing of the $(\text{H}_{2}\text{O})_{2}^{+}$ channel. A and B display transient signals of $(\text{H}_{2}\text{O})_{2}^{+}$ and $\text{H}^{+}$, respectively. C shows the delay- and energy-resolved $\text{H}^{+}$ signals after subtraction of the pump-only contributions. D and E show transient $\text{H}^{+}$ signals for selected kinetic energy ranges. All fits (red lines) were obtained using the parametric model described in Materials and Methods. The data (gray points) in the range (500,50)(-500,-50) fs were excluded from the fitting procedure due to probe-induced pre-excitation effects.
We recorded transient ion-momentum signals capturing the evolution of the $(\text{H}_{2}\text{O})_{2}^{+}$ reaction pathways, illustrated by selected channels of $\text{H}_{3}\text{O}^{+}$ and $\text{H}_{2}\text{O}^{+}$ in Figure 2 A, B and of $(\text{H}_{2}\text{O})_{2}^{+}$ and $\text{H}^{+}$ in Figure 3 A, B. In the inset of Figure 2 A, we show an exemplary velocity-map image of $\text{H}_{3}\text{O}^{+}$; the transient signals were extracted from the central regions (yellow circle) using the three-dimensional momentum distributions, thus excluding contributions from $(\text{H}_{2}\text{O})_{2}^{2+}$. The strong cross-correlation features peaking at time zero are due to overlap of the pump and probe pulses and not discussed further. All ion species display clear time-dependent behavior at positive pump-probe delays $\Delta{t}=t_{\text{probe}}-t_{\text{pump}}>0$. In particular, the $(\text{H}_{2}\text{O})_{2}^{+}$ and $\text{H}_{3}\text{O}^{+}$ transient signals in Figure 2 A and Figure 3 A, respectively, show a time-dependent bleaching, while all other detected channels exhibit an enhancement of the signal relative to the pump-only level. The total ion yield across all signals, outside the cross-correlation region, remains constant. The DP field perturbs the ongoing fragmentation dynamics in $(\text{H}_{2}\text{O})_{2}^{+}$, resulting in a redistribution of the ion signals across different channels. This redistribution can be tracked by analyzing correlations between the transient signals of individual ion species. Comparing the transient signals with one another reveals that the disruption of $\text{H}_{3}\text{O}^{+}$ formation enhances the $\text{H}_{2}\text{O}^{+}$, $\text{OH}^{+}$, and $\text{O}^{+}$ channels, while the bleached $(\text{H}_{2}\text{O})_{2}^{+}$ signal is transferred to the $\text{H}^{+}$ channel.
All acquired pump–probe signals were fitted using a parameterized model based on a linear combination of exponentially-modified-Gaussian (EMG) functions shown as red traces in Figure 2 and 3. Model details and extracted parameters for all measured transient signals are given in Materials and Methods and Supplementary Text. While the $(\text{H}_{2}\text{O})_{2}^{+}$, $\text{H}^{+}$, and $\text{O}^{+}$ traces are well described by a single-term EMG model, the other ion species, $\text{H}_{3}\text{O}^{+}$, $\text{H}_{2}\text{O}^{+}$, and $\text{OH}^{+}$, require a double-term EMG model. Three distinct timescales emerge in the acquired data: The first one represents an ultrafast process of 25…40 fs, corresponding to the drop in the $\text{H}_{3}\text{O}^{+}$ trace. A second timescale of several-hundred femtoseconds corresponds to the recovery of the $\text{H}_{3}\text{O}^{+}$ signal, i. e., 200…250 fs for $\text{H}_{3}\text{O}^{+}$, $\text{H}_{2}\text{O}^{+}$, and $\text{OH}^{+}$ and $\mathord{\sim}580$ fs for the $\text{O}^{+}$ transient. The third timescale, in the $(\text{H}_{2}\text{O})_{2}^{+}$ and $\text{H}^{+}$ signals, is on the order of 1 ps. The identification of the bleached $\text{H}_{3}\text{O}^{+}$ and $(\text{H}_{2}\text{O})_{2}^{+}$ channels as the sources of the observed transients ensures that the extracted timescales correspond directly to the $\text{H}_{3}\text{O}^{+}+\text{OH}$ fragmentation and $(\text{H}_{2}\text{O})_{2}^{+}$ stabilization pathways. Thus, the PT and fragmentation dynamics leading to $\text{H}_{3}\text{O}^{+}$ formation are characterized by lifetimes of $39\pm 9$ fs and $250\pm 28$ fs. The $(\text{H}_{2}\text{O})_{2}^{+}$ stabilization is largely completed at $\mathord{\sim}$1 ps.
The experimental observables were interpreted using ab initio calculations using nonadiabatic surface-hopping dynamics at the extended multistate complete-active-space second-order perturbation (XMS-CASPT2) theory level [2]. The simulations, initialized in the $\text{D}_{0}\ldots\text{D}_{3}$ states and extended to 1 ps, provided time-dependent branching ratios of the dominant channels. Figure 2 C illustrates the total yield of $\text{H}_{3}\text{O}{}\cdots\text{OH}]^{+}$ and $\text{H}_{3}\text{O}^{+}$ products through PT and fragmentation in $\text{D}_{0}$. Single-exponential fits (red traces) to the $\text{H}_{3}\text{O}{}\cdots\text{OH}]^{+}$ formation and $\text{H}_{3}\text{O}^{+}$ fragmentation channels yield the PT and fragmentation timescales. Overall, we obtained excellent agreement of the simulated PT ($38\pm 1$ fs) and fragmentation ($259\pm 19$ fs) timescales with the corresponding experimental timescales derived from the total $\text{H}_{3}\text{O}^{+}$ signal. In contrast, for the simulated dynamics starting in the excited states $\text{D}_{1}\ldots\text{D}_{3}$, the timescales for PT and fragmentation differ markedly from the experimental ones: For PT the simulations yield $70\ldots 85$ fs and for fragmentation they are either considerably slower ($\text{D}_{1}$ & $\text{D}_{2}$, $\mathord{\sim}400$ & $\mathord{\sim}725$ fs) or faster ($\text{D}_{3}$, $\mathord{\sim}200$ fs). This is an intriguing finding, as it suggests the selectivity of the DP technique for our experimental conditions. This selectivity might arise from a resonant transition between the $\text{D}_{0}$ and $\text{D}_{1}$ states, whose ionization-energy difference closely matches the single-photon energy at 800 nm ($\mathord{\sim}1.5$ eV) as indicated in Figure 1 middle panel.
Energy-resolved chemistry
A kinetic-energy-resolved analysis of the $\text{H}_{3}\text{O}^{+}$-formation dynamics yields continuously varying timescales for PT and fragmentation dynamics. The $\text{H}_{3}\text{O}^{+}$ data were divided in ten kinetic-energy ranges from 0 to 1.36 eV, see Figure 2 D, to which we individually fitted the parametric EMG model. These fits were used for background correction of the recorded energy-resolved data. The resulting delay- and kinetic-energy–resolved $\text{H}_{3}\text{O}^{+}$ signals are shown in Figure 2 D. A negative signal corresponds to the probe-induced transfer of $\text{H}_{3}\text{O}^{+}$ population to other fragmentation channels. These signals illustrate the energy-dependence of the chemical dynamics, PT and fragmentation leading to $\text{H}_{3}\text{O}^{+}$. The $\text{H}_{3}\text{O}^{+}$ kinetic-energy redistribution induced by the probe is obtained through integration of these signals along the pump-probe-delay axis. It shows a vast contribution at kinetic energies $\mathord{<}0.3$ eV, see Figure 2 E (blue). This probe-induced kinetic-energy distribution (KED) is clearly different from the observed pump-only $\text{H}_{3}\text{O}^{+}$ KED (yellow) and in excellent agreement with the $\text{H}_{3}\text{O}^{+}$ KED obtained from simulations for ionization to the $\text{D}_{0}$ state of $(\text{H}_{2}\text{O})_{2}^{+}$ (red) [21]. This supports our assignment of observing chemical dynamics in the $\text{D}_{0}$ state, vide supra.
Figure 4: The observed $(\text{H}_{2}\text{O})_{2}^{+}$ dynamics. Energy-resolved PT and fragmentation pathways are identified from the $\text{H}_{3}\text{O}^{+}$ signal and low-energy stabilization via inter-to-intra-molecular vibrational relaxation (IVR) is observed as the $(\text{H}_{2}\text{O})_{2}^{+}$ signal.
An inspection of Figure 2 D reveals distinct timescales associated with the individual kinetic-energy ranges. The timescales extracted for $0\ldots 0.3$ eV are summarized in Figure 2 F. This shows an increase of the fast-component timescale from 19 fs to 60 fs, associated with PT (orange), with a $\text{H}_{3}\text{O}^{+}$ kinetic-energy increase from 0 eV up to 0.15 eV. This slowing demonstrates the transition from a direct to a hindered PT process, indicating changes in the energy flow on the potential energy landscape. This is accompanied by a simultaneous decrease of the slow-component, fragmentation, timescale from 358 fs to 211 fs, associated with faster fragmentation at higher energies (blue). At higher kinetic energies, between 0.15 eV and 0.3 eV, the two timescale converge to $\mathord{\sim}100$ fs, obtained from a modified model assuming equal PT and fragmentation timescales. Thus, strikingly, the $\text{H}_{3}\text{O}^{+}$ formation dynamics exhibits a strong kinetic-energy dependence, with its characteristic timescale changing markedly across the released-energy range, see Figure 4 .
[…]
Bound-state dynamics
We also observed a signal bleach for the $(\text{H}_{2}\text{O})_{2}^{+}$ transient signal shown in Figure 3 A. The parent ion $(\text{H}_{2}\text{O})_{2}^{+}$ does not have kinetic energy from fragmentation, preventing an kinetic-energy–resolved analysis. Nevertheless, the DP induced the transfer of $(\text{H}_{2}\text{O})_{2}^{+}$ population into the $\text{H}^{+}$ channel, Figure 3 B. These $\text{H}^{+}$ signals were analyzed in an energy-resolved manner analogous to the $\text{H}_{3}\text{O}^{+}$ case, providing the energy-resolved signals of $\text{H}^{+}$ shown in Figure 3 C. The time-dependent signals for the energy ranges $0\ldots 0.1$ eV and $1.0\ldots 1.5$ eV are shown in Figure 3 E. The $\text{H}^{+}$ signal in the $0\ldots 0.1$ eV range shows a rapid signal bleach in the first 100 fs, the origin of which remains unresolved. Most $\text{H}^{+}$ appear with kinetic energies of 0.3…1.5 eV, with a timescale of 0.70.71.41.4 ps. We ascribe the $\text{H}^{+}$ signal to the disruption of transient, non-stabilized $(\text{H}_{2}\text{O})_{2}^{+}$ and thus the observed picosecond timescale to $(\text{H}_{2}\text{O})_{2}^{+}$ stabilization, as illustrated in Figure 4 .
However, in the isolated H2O and OH radicals, O-H covalent bonds have dissociation energies of about 5.1 eV and 4.4 eV, respectively [28, 29]. Cleaving such bonds would require the absorption of three to four 800 nm photons. The O-H bonds in isolated $\text{H}_{3}\text{O}^{+}$ are even stronger [30]. Even though the hydronium-ion-hydroxyl-radical pair, like Zundel complexes, has a strong proton delocalization along its O-O axis [31], the proton is overall still strongly bound. Consequently, our weak probing pulse cannot drive the stabilized ion-radical $(\text{H}_{2}\text{O})_{2}^{+}$ pair into the proton channel.
We assume that the weak probe pulse can eject the proton only during the early stage of the dynamics, before the reaction energy is redistributed from intermolecular modes, e. g., proton-shuttling and O-O stretch, to intramolecular modes to form a stabilized $(\text{H}_{2}\text{O})_{2}^{+}$. In other words, the dynamics observed on the hundreds-of-femtoseconds to picosecond timescale are attributed to the decay of transient Zundel-like $(\text{H}_{2}\text{O})_{2}^{+}$, ultimately leading to the formation of a stabilized ion–radical pair $\text{H}_{3}\text{O}^{+}\cdots\text{OH}$. In the liquid, Zundel-like structures decay substantially faster (100–200 fs) [32, 24]. Nevertheless, these sub–200 fs components mainly represent localized proton fluctuations caused by short-lived structural rearrangements, rather than a stable, fully formed Zundel configuration, which persists for much longer times ($\mathord{>}480$ fs) [33].
A remaining open question concerns the electronic-state contribution to the observed stabilization dynamics of $(\text{H}_{2}\text{O})_{2}^{+}$. Based on Ammosov-Delone-Krainov tunneling-ionization calculations [21], the majority of the $(\text{H}_{2}\text{O})_{2}^{+}$ signal detected in the pump-only experiment [17] originates from the $\text{D}_{0}$ and $\text{D}_{1}$ states. However, unlike the assignment of the $\text{H}_{3}\text{O}^{+}$ dynamics to the $\text{D}_{0}$ state (vide supra), the $(\text{H}_{2}\text{O})_{2}^{+}$ signals cannot be uniquely attributed to a single electronic state. A definitive assignment requires further analysis beyond the scope of the present work.
Overall, we unraveled the initial processes after ionization of water exploiting the $(\text{H}_{2}\text{O})_{2}^{+}$ system, with a focus on the appearance of the primary ionic products, including hydronium ion ($\text{H}_{3}\text{O}^{+}$) and stabilized $(\text{H}_{2}\text{O})_{2}^{+}$. Utilizing the energy distributions of the reaction products, our results go beyond electronic-state–averaged descriptions. As illustrated in Figure 4 , low-energy stabilization of $(\text{H}_{2}\text{O})_{2}^{+}$ on picosecond timescales proceeds through Zundel-like configurations that stabilize by atomic rearrangement and vibrational relaxation from intermolecular to intramolecular modes. For increasing energies, we see an energy-driven evolution from sequential ultrafast proton transfer with subsequent fragmentation of the ion-radical pair to an intermediate-timescale concerted pathway.
Our results demonstrate that combining species-controlled samples, multi-mass velocity-map imaging, and disruptive probing with computational-dynamics simulations allows us to disentangle the ultrafast dynamics of complex molecular systems in energy and time. The agreement of our energy-averaged data with earlier ultrafast post-ionization dynamics observed in clusters [2] and liquids [3] demonstrates that our observations are directly relevant to initial-stage liquid-phase chemistry. The sensitivity to internal dynamics, such as the inferred Zundel-like structures, provides opportunities for the investigation of internal motions including general isomerizations. Our approach further opens important prospects for probing biomolecules in their native aqueous environment [34]. Coincidence measurements combining the ion imaging with electron spectroscopy offer direct access to the coupled dynamics of nuclear and electronic reaction pathways. This provides a basis for future attempts to tailor wavepackets to steer ion–radical chemistry in aqueous environments.
Acknowledgments
Funding:
This work was supported by Deutsches Elektronen-Synchrotron (DESY), a member of the Helmholtz Association (HGF), and by the Cluster of Excellence “Advanced Imaging of Matter” of the Deutsche Forschungsgemeinschaft (DFG, AIM, EXC 2056, ID 390715994). Data processing and analysis were carried out using the Maxwell computing resources at DESY. The endstation for controlled-molecule experiments (eCOMO) was funded with support from the Center for Molecular Water Science. I.S.V. acknowledges support from the Alexander von Humboldt Foundation. M.B. and P.S. thank to Czech Science foundation (projects numbers 23-07066S and 26-22810S).
Author contributions:
Original concept: I.S.V., J.K.; Performed experiments: I.S.V., L.B.; Analysis of data: I.S.V. Trajectory based simulations: M.B., P.S.; Software and experimental support: H.B.; Original draft: I.S.V.; Discussion: I.S.V., S.T., M.B., P.S., J.K.; Writing and editing: I.S.V., S.T., P.S., J.K.;
Competing interests:
There are no competing interests to declare.
Data and materials availability:
All data underlying the figures and all analysis scripts are provided as Supplementary Material. All processed data are publicly available at Zenodo (DOI: 10.5281/zenodo.18678891, link). All raw data are archived in DESY’s data repositories under ID 11019141 and are available upon reasonable request.
Supplementary materials
Materials and Methods
Supplementary Text
Tutorial on disruptive probing
Figures SI1 to SI31
Tables SI1 to SI4
References (1-27)

### v2

@@
-URL: ]https://photox.vscht.cz/URL: ]https://www.controlled-molecule-imaging.org
+arXiv:2602.17505v2[physics.chem-ph] 02 Mar 2026
 Abstract
 Aqueous radiation chemistry emerges through ultrafast proton transfer and ion-radical formation with unexplored energy-redistribution dynamics steering the subsequent reactions.
-We performed time-resolved disruptive probing on pure water dimer, $(\text{H}_{2}\text{O})_{2}$, to disentangle the post-ionization reactions.
+We performed time-resolved disruptive-probing experiment of pure water dimer, $(\text{H}_{2}\text{O})_{2}$, to disentangle the post-ionization reactions.
 Through kinetic-energy-resolved ion imaging, we unraveled the dynamics in the $(\text{H}_{2}\text{O})_{2}^{+}$ ground state: at low-energy ($\mathord{\sim}$0.05 eV) ultrafast proton transfer ($\mathord{\sim}$19 fs) is followed by $\text{H}_{3}\text{O}^{+}+\text{OH}$ fragmentation ($\mathord{\sim}$360 fs).
@@
 This reveals how ion-radical formation in ionized hydrogen-bonded networks shapes reactivity in aqueous dynamics.
-Radiation chemistry in aqueous environments is largely governed by water radiolysis [1, 2, 3].
-Upon irradiation, water undergoes a cascade of reactions initiated by the formation of highly reactive ion–radical species [3, 4] that give rise to fundamental phenomena such as biological radiation damage, catalytic transformations, material corrosion, and radical-driven oxidation.
-A molecular-level characterization of transient reactive species formation and the associated energy redistribution is therefore essential for advancing technologies including radiation therapy [4, 5], wastewater treatment [6, 7], nuclear reactor coolant systems [4, 8], and spaceflight applications [9, 10].
+Radiation chemistry in aqueous environments is largely governed by water radiolysis [1, 2].
+Upon irradiation, water undergoes reaction cascades initiated by the formation of highly reactive ion–radical species [3, 4], which underpin biological radiation damage, catalysis, material degradation, and radical-driven oxidation [2].
+A molecular-level characterization of the formation of transient reactive species and the associated energy redistribution is essential for advancing technologies including radiation therapy [2, 5], wastewater treatment [6, 7], nuclear–reactor-coolant systems [2, 8], and spaceflight applications [9, 10].
 Surprisingly, the initial elementary steps driving the subsequent post-ionization dynamics in water were only recently explored in a time-resolved fashion [3, 4, 11].
@@
 A and B display transient signals of the $\text{H}_{3}\text{O}^{+}$ and $\text{H}_{2}\text{O}^{+}$ channels (blue dots and black error bars), respectively.
-Fits (red lines) were obtained using the parametric model described in Materials and Methods.
+Fits (red lines) were obtained using the parametric model described in Materials and Methods SI1.
 The data (gray points) in the range (500,50)(-500,-50) fs were excluded from the fitting procedure due to probe-induced pre-excitation effects.
 The inset in A shows a velocity map of $\text{H}_{3}\text{O}^{+}$ with only the central part (yellow circle) originating from $(\text{H}_{2}\text{O})_{2}^{+}$ dynamics.
-C displays the result of simulations of $(\text{H}_{2}\text{O})_{2}^{+}$ dynamics in its ground electronic state ($\text{D}_{0}$), corresponding data for $\text{D}_{1}$$\text{D}_{3}$ are provided in Supplementary Text.
+C displays the result of simulations of $(\text{H}_{2}\text{O})_{2}^{+}$ dynamics in its ground electronic state ($\text{D}_{0}$), corresponding data for $\text{D}_{1}$$\text{D}_{3}$ are provided in Supplementary Text SI2.
 D represents the background corrected delay- and energy-resolved $\text{H}_{3}\text{O}^{+}$ signal.
 E shows the comparison of the bleached (blue), pump-only (yellow), and simulated ($\text{D}_{0}$, red) kinetic energy distributions [21].
-F displays the timescales of the $\text{H}_{3}\text{O}^{+}$ formation obtained from the kinetic energy-resolved analysis using the parametric models described in Materials and Methods; see also Supplementary Text for the individual fits.
+F displays the timescales of the $\text{H}_{3}\text{O}^{+}$ formation obtained from the kinetic energy-resolved analysis using the parametric models described in Materials and Methods SI1; see also Supplementary Text SI2 for the individual fits.
 Figure 3: Recorded transient signals related to disruptive probing of the $(\text{H}_{2}\text{O})_{2}^{+}$ channel.
@@
 D and E show transient $\text{H}^{+}$ signals for selected kinetic energy ranges.
-All fits (red lines) were obtained using the parametric model described in Materials and Methods.
+All fits (red lines) were obtained using the parametric model described in Materials and Methods SI1.
 The data (gray points) in the range (500,50)(-500,-50) fs were excluded from the fitting procedure due to probe-induced pre-excitation effects.
@@
 All acquired pump–probe signals were fitted using a parameterized model based on a linear combination of exponentially-modified-Gaussian (EMG) functions shown as red traces in Figure 2 and 3.
-Model details and extracted parameters for all measured transient signals are given in Materials and Methods and Supplementary Text.
+Model details and extracted parameters for all measured transient signals are given in Materials and Methods SI1 and Supplementary Text SI2.
 While the $(\text{H}_{2}\text{O})_{2}^{+}$, $\text{H}^{+}$, and $\text{O}^{+}$ traces are well described by a single-term EMG model, the other ion species, $\text{H}_{3}\text{O}^{+}$, $\text{H}_{2}\text{O}^{+}$, and $\text{OH}^{+}$, require a double-term EMG model.
@@
 This is accompanied by a simultaneous decrease of the slow-component, fragmentation, timescale from 358 fs to 211 fs, associated with faster fragmentation at higher energies (blue).
-At higher kinetic energies, between 0.15 eV and 0.3 eV, the two timescale converge to $\mathord{\sim}100$ fs, obtained from a modified model assuming equal PT and fragmentation timescales.
+At higher kinetic energies, between 0.15 eV and 0.3 eV, the two timescales converge to $\mathord{\sim}100$ fs, obtained from a modified model assuming equal PT and fragmentation timescales.
 Thus, strikingly, the $\text{H}_{3}\text{O}^{+}$ formation dynamics exhibits a strong kinetic-energy dependence, with its characteristic timescale changing markedly across the released-energy range, see Figure 4 .
@@
 We also observed a signal bleach for the $(\text{H}_{2}\text{O})_{2}^{+}$ transient signal shown in Figure 3 A.
-The parent ion $(\text{H}_{2}\text{O})_{2}^{+}$ does not have kinetic energy from fragmentation, preventing an kinetic-energy–resolved analysis.
+The parent ion $(\text{H}_{2}\text{O})_{2}^{+}$ does not have kinetic energy from fragmentation, preventing a kinetic-energy–resolved analysis.
 Nevertheless, the DP induced the transfer of $(\text{H}_{2}\text{O})_{2}^{+}$ population into the $\text{H}^{+}$ channel, Figure 3 B.
@@
 All data underlying the figures and all analysis scripts are provided as Supplementary Material.
-All processed data are publicly available at Zenodo (DOI: 10.5281/zenodo.18678891, link).
+All processed data are publicly available at Zenodo (DOI: 10.5281/zenodo.18678891, preview link).
 All raw data are archived in DESY’s data repositories under ID 11019141 and are available upon reasonable request.
 Supplementary materials
-Materials and Methods
-Supplementary Text
-Tutorial on disruptive probing
+Materials and Methods SI1
+Supplementary Text SI2
+Tutorial on disruptive probing SI3
 Figures SI1 to SI31